OpenStudy (anonymous):

How would I solve this? Use the following to answer the question. Triangle a = 12, b = 15, c = 21. The value of s in the formula for area is?

OpenStudy (anonymous):

s is the semi-perimeter. This is called Heron's formula. http://en.wikipedia.org/wiki/Heron's_formula

OpenStudy (anonymous):

\[A = \sqrt{(s)(s-a)(s-b)(s-c)}\]

OpenStudy (anonymous):

s = (a + b + c)/2

OpenStudy (anonymous):

24! Thanks!

OpenStudy (anonymous):

s = 24 A = sqrt(24 * 12 * 9 * 3) >>> a = 24*12*9*3 >>> a 7776 >>> a**0.5 88.18163074019441